OpenStudy (jaylelile):

I'm having a hard time with this problem. 3(x-5)+1=x+4 solve for x

OpenStudy (jaylelile):

@Nnesha ?

OpenStudy (anonymous):

first multiply it out: 3x-15+1=x+4 easier?

OpenStudy (mathmale):

Suggest that you immediately multiply out 3(x-5). Next, combine like terms. Next, isolate all x terms on the left side and all constants on the right side.

OpenStudy (anonymous):

http://www.cymath.com/answer.php?q=3(x-5)%2B1%3Dx%2B4 this website will help

Nnesha (nnesha):

distributive property \[\large\rm \color{ReD}{a}(b+c)=\color{ReD}{a}*b+\color{Red}{a}*c=\color{ReD}{a}b+\color{ReD}{a}c\] multiply both terms inside the parentheses by outside term

OpenStudy (jaylelile):

I got -6.... is this correct?

OpenStudy (mathmale):

You can check your own answer. Merely substitute x =-6 into the orginal equation. If the equation proves to be true, you're right . If not true, you're not.